The New York Ferry (m = 13000.0 kg loaded) is travelling at 11 m/s when the engines are put in reverse. The engine produces a force of 1.0 x 106 N for a period of 20.0 seconds.
1
Expert's answer
2019-03-07T11:24:49-0500

Too much force for so light ferry. Assume its mass is 13 000 000 kg. With this information we can calculate for example:

  • The magnitude of the momentum from the engine:
"\Delta p=F\Delta t=2\cdot 10^7 \text{ kg}\cdot\text{m/s}."


  • The new momentum of the ferry:
"p_1=p_0-\Delta p=mv_0-\Delta p=""=13000000\cdot 11-2\cdot 10^7=123000000\text{ kg}\cdot\text{m/s}."


  • The new velocity of the ferry:
"v_1=p_1/m=9.46\text{ m/s}."